Hi, I need to solve for "F" in the following equation:
P = L + (SQ)/F
the first step I did to isolate F was to move the L,
P-L = (SQ)/F
then do I multiply both sides by 1/SQ to clear the fractions?
Thank you for your help.
Lynn

Answer by stanbon(75887)

solve for "F" in the following equation:
P = L + (SQ)/F
the first step I did to isolate F was to move the L,
P-L = (SQ)/F
---
Invert both sides to get:
F/(SQ) = 1/(P-L)
---
Multiply by SQ to get:
F = (SQ)/(P-L)
===================
Cheers,
Stan H.